The earth's magnetic moment is 8 x 1022 A-m2. If that moment were created by a loop of wire going around the earth (r = 6378 km), what current would need to go through the wire in A?

The wire loop would be on the equator. True or false? Why?

Whether on the equator or close to the equator, which direction would the current travel? Why?

North Pole Lalli till North Pole Magnetic Field Lines EQUATOR Magnetic
Answer #1

The magnetic moment by a loop of wire is given by

Moment = I*A where I is current and A is area

hence 8 x 10^22 = I * pi * 6378000^2 = 19.67 x 10 ^ 8 A

Since the magnetic field are coming out vertically hence the loop of wire should be horizontally so as to match the magnetic field, Also since the magnetic poles are shift a bit hence the loop will be a bit tilted hence the loop will be not exactly at equator.

Since the north pole is downwards hence according to right hand curl rule, the direction of current in loop should be clockwise